A stone is dropped into a well. The sound of the splash is heard 4.80 s later. What is the depth of the well? (Take the speed of sound to be 343 m/s.)
Answer #1

V = 343 m/s The total time taken by the stone to strike the well and the time taken by the sound to come up is, t + t, = 4.818t1 =v. (4.8 s-1) Substitute the values in above equation. n/s) t =(343 m/s) (4.8 s-4) (4.9 m/s> ) t =(343 m/s)(4.8 st) t =(